OpenStudy (anonymous):

Two cars leave town, one driving North the other East. They are 104 miles apart when one of them is 56 miles farther from town than the other. At that time how far were they from each town?

OpenStudy (amistre64):

x^2 + y^2 = r^2 x = y+56 (y+56)^2 + y^2 = r^2

OpenStudy (amistre64):

r = 104

OpenStudy (amistre64):

(y+56)^2 + y^2 = 104^2 solve for y :)

OpenStudy (anonymous):

how do i solve for y with all of those squares